using quadratic formula
4x^2-12x+11=0

Answer by stanbon(75887)

4x^2-12x+11=0
x = [12 +- sqrt(144-4*4*11)]/8
x = [12 +- sqrt(-32)]/8
x = [12 +- 4sqrt(2)i]/8
x = (3/2) + (1/2)sqrt(2)i or x = (3/2) - (1/2)sqrt(2)i
=======================
Cheers,
Stan H.